OpenStudy (anonymous):

How do I factor (1331x^3)-8

OpenStudy (anonymous):

1331^3 - 8 (11^3)^3 - 8 11 ^9 - 8

OpenStudy (anonymous):

1331 = 11 ^ 3 1331 ^ 3 = 11 (3 x 3) = 11 ^ 9

OpenStudy (anonymous):

Sorry 1331 ^ 3 = 11 ^ (3 x 3) = 11 ^ 9

OpenStudy (anonymous):

\[1331X ^{3}-8\]

OpenStudy (anonymous):

this the equation.

OpenStudy (anonymous):

You can't eliminate the X.

OpenStudy (anonymous):

Sorry (11x) ^3 - 8

OpenStudy (anonymous):

it is ok

OpenStudy (anonymous):

Just to be clear: = (11x)³ - 2³

OpenStudy (anonymous):

Ah good point Chlorophyll ;D